Pittsburgh Custom Products (PCP) purchased a new machine for ram-cambering large I beams. PCP expects to bend 65 beams at $2,200 per beam in each of the first 3 years, after which it expects to bend 100 beams per year at $2,700 per beam through year 10. If the company’s minimum attractive rate of return is 12% per year, what is the present worth of the expected revenue?
Answer #1

Solution:

P = 65 (1200) (P/A, 12%,3) + 100 (2700) (P/A, 12%,7) (P/F, 12%,3)

P = (65 * 1200 * 2.4018) + (100 * 2700 * 4.5638 * 0.7118)

P = 187,340 + 877,098.47

P = 1,064,439

Thus present worth of the expected revenue is $1,064,439